# Towards Agentic Intelligence for Materials Science

### v1

@@ 3.1.2 Classification Tasks
Classification is a fundamental task in materials science, referring to the automatic categorization of materials into predefined classes based on their composition, structure, or other characteristic data 429, 162. Classification is a crucial step in numerous downstream applications, from interpreting experimental data to controlling material quality and accelerating the discovery of new materials 100. Meanwhile, because LLMs can effectively handle noisy experimental and computational material data and integrate data inputs from multiple modalities such as text, images, and numerical data 207, 310, they are gradually becoming a general solution for classification problems.
@@
The classification of defects—whether vacancies, dislocations, or grain boundaries—is an important aspect of predicting material performance and ensuring safety and optimization of manufacturing yield. In practice, defect signals may be derived from microscopy 274, in-situ sensors 1, or the operational setting ranges from offline labeling to on-the-fly process monitoring 107. LLM has broad application prospects in this field, serving both static analysis tasks such as defect labeling and integration into closed-loop systems to provide real-time decision support for researchers, thereby accelerating research and development cycles 54, 45. At present, most deployments still treat classification as a modular, reactive component; closed-loop autonomy typically requires additional planning, control, and verification modules beyond classification alone.
In additive manufacturing, LLMs fine-tuned on process parameter data can predict defect regimes (i.e. keyholing or lack of fusion) directly from natural language descriptions 271. This enables faster build parameter optimization while significantly reducing reliance on resource-intensive simulations and physical experiments. Similarly, in materials characterization, conversational LLM systems have been developed to support defect classification from Atomic Force Microscopy (AFM) images. These typically involve a two-step pipeline whereby a vision-based model performs an initial defect classification, and then an LLM acts as a conversational assistant. In this latter role, the LLM contextualizes the classification based on queries from users and experimental parameters and provides actionable feedback in natural language, such as suggesting adjustments of scan speed to minimize artifacts 42. More broadly, classification becomes substantially more valuable when embedded in an interactive workflow, but the classification objective itself remains a proxy label unless the surrounding system optimizes end-to-end discovery outcomes 296. This evolution transforms the LLM from a passive classifier into an active participant in the experimental workflow-capable of hierarchical reasoning linking sensor input to corrective measures with reduced reliance on continuous expert supervision 207. Finally, beyond classification, LLM-driven agents are being developed that automate the complete scientific discovery process, from hypothesis generation 316 to experimental design and data analysis 45.
@@ 3.1.3 Advanced Methodologies
While GNNs perform well in modeling local atomic environments, their main drawbacks are related to modeling long-range interactions and a well-known issue of over-smoothing in deep architectures 266. Transformers specialize in the opposite: global dependencies with high computational costs for large structures 364. Hybrid GNN-Transformer architectures seek to harness the strengths of both approaches 409.
These hybrid models generally use a GNN to learn efficiently local structural features from predefined graphs. The node embeddings produced by this step encode structure-aware information and are treated as a sequence of tokens by a Transformer module. This latter component then uses self-attention to infer long-range interactions within the material 99. The hierarchical method therefore overcomes the individual limitations of both architectures by leveraging their combined strengths to afford even more accurate and scalable predictions of material properties 234, 216. This synergy thus allows for a rich representation of materials from the atomic to the macroscopic scale. However, stronger representations alone do not guarantee improved out-of-distribution reliability, motivating uncertainty-aware evaluation and closed-loop validation.
@@
The physical interdependencies between key material properties (such as band gap, elastic modulus, and formation energy) make MTL a very suitable strategy for building generalizable models 307. By training a single model on these related tasks simultaneously, MTL seeks to discover a unified feature space that inherently encodes the underlying shared representations consistent with physical correlations, resulting in a more robust and generalizable representation and leading to better prediction performance 60. A practical limitation is negative transfer: performance can degrade when tasks conflict or when task data are imbalanced and heterogeneous, as is common in materials datasets 27.
Architectures such as the Multi-Task Pre-trained Transformer Encoder (MTENCODER) learn general material representations from diverse properties 283. By combining the signals from several auxiliary activities, this method makes data more efficient and helps the models generalize better to new tasks and materials. Thus, using MTL in transformer encoders creates strong and transferable embeddings that are an important step toward flexible foundation models for materials science 321. From a pipeline-centric perspective, MTL performs internal credit sharing across proxy tasks; aligning task selection and weighting with discovery-level objectives remains an open systems challenge 305.
@@
To make AI-driven material discovery effective, we need a system capable of accurately measuring prediction uncertainty. This measurement of uncertainty is very important for building trust in AI systems, directing active learning strategies, and reliably finding samples that are outside of the distribution 351.
In closed-loop discovery, uncertainty is also a decision variable (e.g., acquisition or stopping). There are two main kinds of uncertainty quantification methods: aleatoric uncertainty, which comes from noise in the data, and epistemic uncertainty, arising from limited data coverage and model misspecification 182. There are different ways to figure out prediction intervals for material properties. Some of these are probabilistic methods like Gaussian processes, while others are quantile loss functions and methods that directly model prediction error 351. Having good estimates of epistemic uncertainty is very helpful because they show parts of chemical space where the model is not very certain. This directly benefits targeted data collection, which means that future experiments or simulations could be guided by an active learning loop. This makes the process of discovering new findings faster and uses fewer resources 159. In materials settings, uncertainty estimation is further complicated by multi-fidelity data (simulation vs. experiment), distribution shift across conditions, and the compositional long tail—precisely where naïve confidence estimates can fail. More and more studies are now trying to determine how much uncertainty there is in the outputs of LLMs. This is a necessary step for using them reliably in important scientific work 398. Finally, while uncertainty quantification can make proxy-task predictors more reliable, end-to-end discovery requires propagating uncertainty through multi-step planning and experimentation decisions.
@@ 3.2.1 Information Extraction
Information extraction in materials science focuses on converting experimental descriptions and performance data from publications into structured formats suitable for data-driven research. As a pipeline-critical bottleneck, extraction determines what becomes trainable downstream (datasets, benchmarks, and ultimately the objectives that models optimize). Early works used rule-based text mining systems such as ChemDataExtractor 339, which automatically produced databases of synthesis parameters 185, battery compositions 154, thermoelectric properties 322, and mechanical data 197 from journal article. These approaches provided reproducible extraction pipelines but were constrained by fixed grammars and limited adaptability to variations in scientific writing.
To address these limitations, subsequent research introduced statistical and representation-learning approaches. Embedding models trained on large-scale materials corpora captured latent chemical and physical relations through co-occurrence patterns 358, while domain-specific advances in named entity recognition and normalization improved terminology consistency 389. Building on these foundations, transformer-based models such as MatSciBERT 145 and MechBERT 198 provided contextualized embeddings that greatly improved extraction robustness and accuracy across diverse publication sources.
Research on LLMs for material science, utilizing prompt engineering and fine-tuning, has expanded the limits of extraction system. ChatExtract 281 builds a conversational system with engineered prompts and follow-up questions to identify evidence sentences, extract structured values, and self-verify, reaching high precision and recall suitable for database curation. Similarly, MaTableGPT 405, an instruction-tuned and table-aware model, jointly reasons over narrative text and tabular data for quantitative extraction. da Silva et al. 88 presented an automated pipeline for extracting synthesis procedures for reticular materials from PDFs using prompt-based paragraph classification. Wang et al. 373 also explored LLM-augmented decision programs that link extracted synthesis data to design workflows. At the same time, others combine extraction and verification mechanisms to autonomously construct structural datasets from full-text literature 144, 126, 310. Further,  Ansari and Moosavi 22 developed a chemist AI agent that extracts and structures materials science data with chain-of-verification tool to create datasets. Collectively, these efforts suggest a shift from one-shot parsing toward agentic extraction pipelines, and yet extraction still functions largely as a reactive module unless it is coupled to discovery-level objectives and outcome-grounded evaluation.
@@ 3.2.2 Knowledge Graph
As extraction quality improved, researchers began to integrate extracted entities and relations into structured knowledge representations. Hybrid and semi-supervised methods complement purely model-based extraction by building structured knowledge representations. Hybrid methods combine GNNs, structured ontology models and LLMs to extend knowledge graphs and interpret predicted relations 158. Semi-supervised pipelines have been applied to identify synthesis–processing–property relations in alloy systems 376. Machine learning combined with text mining has enabled the reconstruction of magnetic and superconducting phase diagrams from literature 86. These extracted relations feed into knowledge graphs such as MatKG and terminology-based networks 366, 421, which integrate entities, parameters, and outcomes into a structured representation for downstream reasoning across the composition–structure–property space. Prein et al. 284 proposed a ranking-based approach to inorganic synthesis planning that learns precursor relationships in a latent space, providing new insights for reaction network construction. A key limitation is that KG coverage and correctness remain bounded by extraction noise and ontology choices, and thus errors could propagate into downstream reasoning and planning 276.
@@ 3.2.3 Database Automation and Annotation
Large automatically generated corpora 402 enabled standardized benchmarks for entity and relation extraction. Together, these efforts shifted the extraction paradigm from rigid rules to learned semantic representations. Frameworks like MatNexus 413 and MatScIE 141 provide modular pipelines to aggregate extraction outputs and align them with ontologies. CALMS 285 proposed linking LLM-based extraction with instrument control and database queries. Multimodal extraction combining text and tables 414 and literature-based reconstruction of the materials tetrahedron 148 illustrated progress toward unified modeling of composition, structure, process, and property. From a pipeline-centric perspective, automated corpora and benchmarks can accelerate iteration, but they can also freeze proxy objectives when benchmark definitions drift from discovery outcome-grounded evaluation and continual dataset expansion and revision. In practice, rule-free and model-agnostic curation workflows now turn full texts into mid-sized, high-precision databases with minimal bespoke coding and near-perfect precision–recall trade-offs, making them ideal for rapid property-focused datasets.
@@ 3.3.1 Structure Generation
The task of structure generation in AI4Materials spans multiple scales and material classes, aiming to propose new candidate structures that are valid, stable, and potentially synthesizable 410. In practice, these criteria are often operationalized using surrogates (e.g., predicted formation energy), which can diverge from real synthesis feasibility—motivating the pipeline-centric emphasis on verification and closed-loop feedback. While early research focused on crystalline solids—where periodicity, symmetry, and lattice constraints dominate—the field now extends to polymers, porous frameworks (MOFs 249/COFs 184), amorphous phases such as glasses and electrolytes, microstructures at the mesoscale, and even architected metamaterials. In these fields, materials can be represented as matrices, voxels, or graphs, and generative models 243 are used to generate realistic atomic structures. For discovery, the key differentiators are less the model family than whether generation is coupled to evaluation in the loop.
For crystalline materials, a complementary research approach has emerged that represents crystal structures as text sequences and generates them using language models. This language-based paradigm treats structural information as a sequence of tokens, which makes it possible to generate lattices and compositions in an autoregressive manner. Following this idea, CrystalLLM26 further highlighted the potential of sequential generation by applying autoregressive language modeling to produce CIF file sequences directly. Gruver et al. 138 further demonstrated that fine-tuned LLaMA-2 models can directly generate stable inorganic crystal structures purely as text, while achieving nearly double the generation speed that of diffusion-based baselines (i.e. CDVAE392). More recently, MATLLMSEARCH 116 showed that pre-trained LLMs can act as agentic generators of crystal structures without additional fine-tuning. Their system autonomously explores and optimizes materials by combining a pre-trained LLaMA-3.1 model with evolutionary search and physical evaluation. This shows a shift in regarding LLMs as reasoning agents that can iteratively search, evaluate, and improve candidate structures. Collectively, these works mark a transition from generative sampling to search-with-verification, which is loser to our pipeline-centric view. Unfortunately, most evaluations still rely primarily on simulated or surrogate validators.
@@ 3.3.2 Inverse Design
In materials generation using generative models, the objective is typically to produce new materials that follow the distribution of the training data. However, in practical materials discovery, there is often a demand to identify materials that exhibit specific desired properties. MatAgent 344, proposes an agentic material generation framework combining an LLM, a diffusion model, and a property prediction model. In this framework, the LLM proposes candidate compositions and iteratively refines them based on feedback from the property predictor, enabling the autonomous generation of materials with desired properties. RL has also entered this space: to inverse inorganic oxide materials design to target promising compounds using specified property and synthesis objectives 272. Because RL naturally supports long-horizon objectives, careful reward and constraint design is essential to avoid optimizing proxy signals that do not translate to real discovery outcomes. Deep dreaming solves the inverse design challenge for metal–organic frameworks by directly optimizing string-based representations of MOF linkers within a unified neural network to generate novel, synthesizable structures with targeted properties 82. MatDesINNe is a framework based on invertible neural networks (INNs) that learns both forward and reverse mappings between design parameters and target properties, enabling generation of candidate materials for specified property goals (e.g., band gaps in 2D materials) 114. PolyTAO (Transformer-Assisted Oriented model) is a generative pretrained model for polymers: given a vector of target polymer properties, it generates polymer repeat unit SMILES with very high chemical validity and accurately matches desired property values in semi-template or even template-free settings 291. These approaches share a common dependency on surrogate property models and feasibility constraints, motivating closed-loop verification later in the discovery pipeline.

### v2

@@ 3.1.2 Classification Tasks
Classification is a fundamental task in materials science, referring to the automatic categorization of materials into predefined classes based on their composition, structure, or other characteristic data 432, 163. Classification is a crucial step in numerous downstream applications, from interpreting experimental data to controlling material quality and accelerating the discovery of new materials 101. Meanwhile, because LLMs can effectively handle noisy experimental and computational materials data and integrate data inputs from multiple modalities such as text, images, and numerical data 208, 312, they are gradually becoming a general solution for classification problems.
@@
The classification of defects—whether vacancies, dislocations, or grain boundaries—is an important aspect of predicting materials performance and ensuring safety and optimization of manufacturing yield. In practice, defect signals may be derived from microscopy 276, in-situ sensors 1, or the operational setting ranges from offline labeling to on-the-fly process monitoring 108. LLM has broad application prospects in this field, serving both static analysis tasks such as defect labeling and integration into closed-loop systems to provide real-time decision support for researchers, thereby accelerating research and development cycles 55, 46. At present, most deployments still treat classification as a modular, reactive component; closed-loop autonomy typically requires additional planning, control, and verification modules beyond classification alone.
In additive manufacturing, LLMs fine-tuned on process parameter data can predict defect regimes (i.e. keyholing or lack of fusion) directly from natural language descriptions 273. This enables faster build parameter optimization while significantly reducing reliance on resource-intensive simulations and physical experiments. Similarly, in materials characterization, conversational LLM systems have been developed to support defect classification from Atomic Force Microscopy (AFM) images. These typically involve a two-step pipeline whereby a vision-based model performs an initial defect classification, and then an LLM acts as a conversational assistant. In this latter role, the LLM contextualizes the classification based on queries from users and experimental parameters and provides actionable feedback in natural language, such as suggesting adjustments of scan speed to minimize artifacts 43. More broadly, classification becomes substantially more valuable when embedded in an interactive workflow, but the classification objective itself remains a proxy label unless the surrounding system optimizes end-to-end discovery outcomes 298. This evolution transforms the LLM from a passive classifier into an active participant in the experimental workflow-capable of hierarchical reasoning linking sensor input to corrective measures with reduced reliance on continuous expert supervision 208. Finally, beyond classification, LLM-driven agents are being developed that automate the complete scientific discovery process, from hypothesis generation 318 to experimental design and data analysis 46.
@@ 3.1.3 Advanced Methodologies
While GNNs perform well in modeling local atomic environments, their main drawbacks are related to modeling long-range interactions and a well-known issue of over-smoothing in deep architectures 268. Transformers specialize in the opposite: global dependencies with high computational costs for large structures 367. Hybrid GNN-Transformer architectures seek to harness the strengths of both approaches 412.
These hybrid models generally use a GNN to learn efficiently local structural features from predefined graphs. The node embeddings produced by this step encode structure-aware information and are treated as a sequence of tokens by a Transformer module. This latter component then uses self-attention to infer long-range interactions within the material 100. The hierarchical method therefore overcomes the individual limitations of both architectures by leveraging their combined strengths to afford even more accurate and scalable predictions of material properties 236, 218. This synergy thus allows for a rich representation of materials from the atomic to the macroscopic scale. However, stronger representations alone do not guarantee improved out-of-distribution reliability, motivating uncertainty-aware evaluation and closed-loop validation.
@@
The physical interdependencies between key material properties (such as band gap, elastic modulus, and formation energy) make MTL a very suitable strategy for building generalizable models 309. By training a single model on these related tasks simultaneously, MTL seeks to discover a unified feature space that inherently encodes the underlying shared representations consistent with physical correlations, resulting in a more robust and generalizable representation and leading to better prediction performance 61. A practical limitation is negative transfer: performance can degrade when tasks conflict or when task data are imbalanced and heterogeneous, as is common in materials datasets 28.
Architectures such as the Multi-Task Pre-trained Transformer Encoder (MTENCODER) learn general material representations from diverse properties 285. By combining the signals from several auxiliary activities, this method makes data more efficient and helps the models generalize better to new tasks and materials. Thus, using MTL in transformer encoders creates strong and transferable embeddings that are an important step toward flexible foundation models for materials science 323. From a pipeline-centric perspective, MTL performs internal credit sharing across proxy tasks; aligning task selection and weighting with discovery-level objectives remains an open systems challenge 307.
@@
To make AI-driven materials discovery effective, we need a system capable of accurately measuring prediction uncertainty. This measurement of uncertainty is very important for building trust in AI systems, directing active learning strategies, and reliably finding samples that are outside of the distribution 353.
In closed-loop discovery, uncertainty is also a decision variable (e.g., acquisition or stopping). There are two main kinds of uncertainty quantification methods: aleatoric uncertainty, which comes from noise in the data, and epistemic uncertainty, arising from limited data coverage and model misspecification 183. There are different ways to figure out prediction intervals for material properties. Some of these are probabilistic methods like Gaussian processes, while others are quantile loss functions and methods that directly model prediction error 353. Having good estimates of epistemic uncertainty is very helpful because they show parts of chemical space where the model is not very certain. This directly benefits targeted data collection, which means that future experiments or simulations could be guided by an active learning loop. This makes the process of discovering new findings faster and uses fewer resources 160. In materials settings, uncertainty estimation is further complicated by multi-fidelity data (simulation vs. experiment), distribution shift across conditions, and the compositional long tail—precisely where naïve confidence estimates can fail. More and more studies are now trying to determine how much uncertainty there is in the outputs of LLMs. This is a necessary step for using them reliably in important scientific work 401. Finally, while uncertainty quantification can make proxy-task predictors more reliable, end-to-end discovery requires propagating uncertainty through multi-step planning and experimentation decisions.
@@ 3.2.1 Information Extraction
Information extraction in materials science focuses on converting experimental descriptions and performance data from publications into structured formats suitable for data-driven research. As a pipeline-critical bottleneck, extraction determines what becomes trainable downstream (datasets, benchmarks, and ultimately the objectives that models optimize). Early works used rule-based text mining systems such as ChemDataExtractor 341, which automatically produced databases of synthesis parameters 186, battery compositions 155, thermoelectric properties 324, and mechanical data 198 from journal article. These approaches provided reproducible extraction pipelines but were constrained by fixed grammars and limited adaptability to variations in scientific writing.
To address these limitations, subsequent research introduced statistical and representation-learning approaches. Embedding models trained on large-scale materials corpora captured latent chemical and physical relations through co-occurrence patterns 361, while domain-specific advances in named entity recognition and normalization improved terminology consistency 392. Building on these foundations, transformer-based models such as MatSciBERT 146 and MechBERT 199 provided contextualized embeddings that greatly improved extraction robustness and accuracy across diverse publication sources.
Research on LLMs for materials science, utilizing prompt engineering and fine-tuning, has expanded the limits of extraction system. ChatExtract 283 builds a conversational system with engineered prompts and follow-up questions to identify evidence sentences, extract structured values, and self-verify, reaching high precision and recall suitable for database curation. Similarly, MaTableGPT 408, an instruction-tuned and table-aware model, jointly reasons over narrative text and tabular data for quantitative extraction. da Silva et al. 89 presented an automated pipeline for extracting synthesis procedures for reticular materials from PDFs using prompt-based paragraph classification. Wang et al. 376 also explored LLM-augmented decision programs that link extracted synthesis data to design workflows. At the same time, others combine extraction and verification mechanisms to autonomously construct structural datasets from full-text literature 145, 127, 312. Further,  Ansari and Moosavi 23 developed a chemist AI agent that extracts and structures materials science data with chain-of-verification tool to create datasets. Collectively, these efforts suggest a shift from one-shot parsing toward agentic extraction pipelines, and yet extraction still functions largely as a reactive module unless it is coupled to discovery-level objectives and outcome-grounded evaluation.
@@ 3.2.2 Knowledge Graph
As extraction quality improved, researchers began to integrate extracted entities and relations into structured knowledge representations. Hybrid and semi-supervised methods complement purely model-based extraction by building structured knowledge representations. Hybrid methods combine GNNs, structured ontology models and LLMs to extend knowledge graphs and interpret predicted relations 159. Semi-supervised pipelines have been applied to identify synthesis–processing–property relations in alloy systems 379. Machine learning combined with text mining has enabled the reconstruction of magnetic and superconducting phase diagrams from literature 87. These extracted relations feed into knowledge graphs such as MatKG and terminology-based networks 369, 424, which integrate entities, parameters, and outcomes into a structured representation for downstream reasoning across the composition–structure–property space. Prein et al. 286 proposed a ranking-based approach to inorganic synthesis planning that learns precursor relationships in a latent space, providing new insights for reaction network construction. A key limitation is that KG coverage and correctness remain bounded by extraction noise and ontology choices, and thus errors could propagate into downstream reasoning and planning 278.
@@ 3.2.3 Database Automation and Annotation
Large automatically generated corpora 405 enabled standardized benchmarks for entity and relation extraction. Together, these efforts shifted the extraction paradigm from rigid rules to learned semantic representations. Frameworks like MatNexus 416 and MatScIE 142 provide modular pipelines to aggregate extraction outputs and align them with ontologies. CALMS 287 proposed linking LLM-based extraction with instrument control and database queries. Multimodal extraction combining text and tables 417 and literature-based reconstruction of the materials tetrahedron 149 illustrated progress toward unified modeling of composition, structure, process, and property. From a pipeline-centric perspective, automated corpora and benchmarks can accelerate iteration, but they can also freeze proxy objectives when benchmark definitions drift from discovery outcome-grounded evaluation and continual dataset expansion and revision. In practice, rule-free and model-agnostic curation workflows now turn full texts into mid-sized, high-precision databases with minimal bespoke coding and near-perfect precision–recall trade-offs, making them ideal for rapid property-focused datasets.
@@ 3.3.1 Structure Generation
The task of structure generation in AI4Materials spans multiple scales and materials classes, aiming to propose new candidate structures that are valid, stable, and potentially synthesizable 413. In practice, these criteria are often operationalized using surrogates (e.g., predicted formation energy), which can diverge from real synthesis feasibility—motivating the pipeline-centric emphasis on verification and closed-loop feedback. While early research focused on crystalline solids—where periodicity, symmetry, and lattice constraints dominate—the field now extends to polymers, porous frameworks (MOFs 251/COFs 185), amorphous phases such as glasses and electrolytes, microstructures at the mesoscale, and even architected metamaterials. In these fields, materials can be represented as matrices, voxels, or graphs, and generative models 245 are used to generate realistic atomic structures. For discovery, the key differentiators are less the model family than whether generation is coupled to evaluation in the loop.
For crystalline materials, a complementary research approach has emerged that represents crystal structures as text sequences and generates them using language models. This language-based paradigm treats structural information as a sequence of tokens, which makes it possible to generate lattices and compositions in an autoregressive manner. Following this idea, CrystalLLM27 further highlighted the potential of sequential generation by applying autoregressive language modeling to produce CIF file sequences directly. Gruver et al. 139 further demonstrated that fine-tuned LLaMA-2 models can directly generate stable inorganic crystal structures purely as text, while achieving nearly double the generation speed that of diffusion-based baselines (i.e. CDVAE395). More recently, MATLLMSEARCH 117 showed that pre-trained LLMs can act as agentic generators of crystal structures without additional fine-tuning. Their system autonomously explores and optimizes materials by combining a pre-trained LLaMA-3.1 model with evolutionary search and physical evaluation. This shows a shift in regarding LLMs as reasoning agents that can iteratively search, evaluate, and improve candidate structures. Collectively, these works mark a transition from generative sampling to search-with-verification, which is loser to our pipeline-centric view. Unfortunately, most evaluations still rely primarily on simulated or surrogate validators.
@@ 3.3.2 Inverse Design
In materials generation using generative models, the objective is typically to produce new materials that follow the distribution of the training data. However, in practical materials discovery, there is often a demand to identify materials that exhibit specific desired properties. MatAgent 346, proposes an agentic materials generation framework combining an LLM, a diffusion model, and a property prediction model. In this framework, the LLM proposes candidate compositions and iteratively refines them based on feedback from the property predictor, enabling the autonomous generation of materials with desired properties. RL has also entered this space: to inverse inorganic oxide materials design to target promising compounds using specified property and synthesis objectives 274. Because RL naturally supports long-horizon objectives, careful reward and constraint design is essential to avoid optimizing proxy signals that do not translate to real discovery outcomes. Deep dreaming solves the inverse design challenge for metal–organic frameworks by directly optimizing string-based representations of MOF linkers within a unified neural network to generate novel, synthesizable structures with targeted properties 83. MatDesINNe is a framework based on invertible neural networks (INNs) that learns both forward and reverse mappings between design parameters and target properties, enabling generation of candidate materials for specified property goals (e.g., band gaps in 2D materials) 115. PolyTAO (Transformer-Assisted Oriented model) is a generative pre-trained model for polymers: given a vector of target polymer properties, it generates polymer repeat unit SMILES with very high chemical validity and accurately matches desired property values in semi-template or even template-free settings 293. These approaches share a common dependency on surrogate property models and feasibility constraints, motivating closed-loop verification later in the discovery pipeline.